OpenStudy (anonymous):

Another one: (tan^2(t))/(sec(t))=sec(t)-cos(t) I could only simplify to (1-cos(t))/(cos(t))=sec(t)-cos(t) But then again that's probably wrong.

OpenStudy (abdullahm):

Hint: tan^2 (t) = sin^2 (t)/ cos^2 (t) sec (t) = 1/cos (t) Plug both of those into the left hand side and right hand side. And simplify. You'll need to simplify a complex fraction on the LHS, and you'll need to get a common denominator on the RHS. And then on the RHS, you use: 1 - cos^2 (t) = sin^2 (t)

OpenStudy (abdullahm):

If you follow all that, you should be able to prove that identity! I have to go, so I tried to summarize all you needed to do to solve your question! I hope that helps. I'll try to come on later to see if you have any more questions. :)

OpenStudy (anonymous):

Thanks I think I got it now:)

OpenStudy (abdullahm):

If you post your work and solution, I can make sure you did it correctly! :)